Calculate the extra compound interest received when R18 400,00 invested at 13,2% per annum for 11,5 years, interest paid annually, is invested at the same rate but with interest paid monthly.


1
Expert's answer
2021-05-28T10:21:59-0400

interest paid annually:

"FV_1=18400(1+0.132)^{11.5}=R\ 76568.78"


 interest paid monthly:

"FV_2=18400(1+0.132/12)^{11.5\cdot12}=R\ 83267.84"


extra compound interest:

"FV_2-FV_1=83267.84-76568.78=R\ 6699.06"